# Learning Linearity in Audio Consistency Autoencoders via Implicit Regularization

### v1

@@ 4.2 Reconstruction Quality
We evaluate the reconstruction quality of the AEs on 2-second audio chunks ($0.5$s overlap) from the MusicCaps dataset [37]. We reconstruct the full track using overlap-add and compute three metrics: (I) The Signal to Noise Ratio (SNR) on the waveform. (II) A Multi-Scale Spectral distance (MSS) between the original and reconstructed log-mel spectrograms using multiple resolutions3338080 mel bands and hop lengths of $\approx$ $10$, $25$, and $50$ ms.. (III) The Kernel Audio Distance (KAD) [38] between the embedding distributions of the original and reconstructed tracks [^4] Embedding distribution metrics have become a standard way of evaluating generative models, so we include it for completeness. KAD is computed using half of the reconstructed tracks, using the other half as the reference set. We use the kadtk library with the LAION-CLAP model..
Table 2 summarizes the results. Our retrained baseline (M2L) already improves on the public M2L-Pub weights [4], likely due to our distinct training data, batch size and augmentations. The linearized model (Lin-CAE) remains notably comparable on MSS and improves further on SNR, indicating that our method does not degrade reconstruction fidelity. Compared to SA-VAE, Lin-CAE is slightly worse, a similar finding to [4, 26]. We note that SA-VAE uses different data and a two-stage training procedure that optimizes both reconstruction and adversarial losses, while Lin-CAE is trained with a single denoising objective. KAD scores for Lin-CAE are better than our retrained M2L-Pub model and SA-VAE , but are worse than the public M2L weights (possibly reflecting the different training data distribution).
@@ 4.4 Additivity
Additivity is evaluated on four source mixtures from the Musdb18-HQ test set [39], where a mixture is : $\text{mix}=\sum^{4}_{i=1}s_{i}$. $s_{i}$ correspond to either vocals, bass, drum or other. We measure Encoder Additivity Error as the deviation of a mixture’s latent from the sum of its source latents, calculated as the relative $L_{2}$ error: $\|\operatorname{Enc}_{\theta}(\text{mix})-\sum_{i}\operatorname{Enc}_{\theta}(s_{i})\|_{2}/\|\operatorname{Enc}_{\theta}(\text{mix})\|_{2}$. For Decoder Additivity (Composability), we test the ability to reconstruct a mixture from the sum of its source latents by SNR and MSS between the decoded sum of latents, $\operatorname{Dec}_{\theta}(\sum_{i}\operatorname{Enc}_{\theta}(s_{i}))$, to the autoencoded mixture, $\operatorname{Dec}_{\theta}(\operatorname{Enc}_{\theta}(\text{mix}))$.
@@ 4.5 Music source separation via latent arithmetic
We perform source separation via latent arithmetic on the Musdb18-HQ test set. A source $\hat{s}_{i}$ is estimated by subtracting the latent representation of its accompaniment from the latent of the full mixture: $\hat{s}_{i}=\operatorname{Dec}_{\theta}(\operatorname{Enc}_{\theta}(\text{mix})-\operatorname{Enc}_{\theta}(\text{mix}-s_{i}))$. Figure 3 illustrates this process. The quality of the separated source is evaluated against the autoencoded ground-truth source, $\operatorname{Dec}_{\theta}(\operatorname{Enc}_{\theta}(s_{i}))$, using Scale-invariant Signal to Distortion Ratio (SI-SDR) [40] and MSS.
Table 1 shows that Lin-CAE significantly outperforms the baselines in latent arithmetic/oracle source separation tasks, achieving higher SI-SDR and lower MSS scores by a large margin for every instrument. Notably, both separation metrics are of the same order of magnitude as the reconstruction values for the mix, which can be seen as an upper bound for the model. That is not the case, however, for any of the baselines, with the least significant gap for SA-VAE. We note that during training, we have only shown positive gains to the model, while during separation, we are effectively applying negative gains when subtracting latents. This further indicates the robustness of the learned linear structure.
@@ 5 Conclusion
We introduce a simple method to induce an approximately linear latent space in audio autoencoders during training. Using implicit conditioning via data augmentation, our approach enforces the properties of a linear map (homogeneity and additivity) without necessitating any modifications to the architecture or loss function. Experiments with Consistency Autoencoders (CAEs) show that linearization preserves high-quality reconstruction and can be used for source separation via latent arithmetic.
Our approach enables more interpretable and controllable audio generation. Future work could extend this method to improved CAE architectures [26] and downstream tasks like generative source separation [41]. Direct, low-level manipulations in compressed space promise new possibilities for audio editing and efficient signal processing.

### v2

@@ 4.2 Reconstruction Quality
We evaluate the reconstruction quality of the AEs on 2-second audio chunks ($0.5$s overlap) from the MusicCaps dataset [agostinelliMusicLMGeneratingMusic2023]. We reconstruct the full track using overlap-add and compute three metrics: (I) The Signal to Noise Ratio (SNR) on the waveform. (II) A Multi-Scale Spectral distance (MSS) between the original and reconstructed log-mel spectrograms [^3] We use mel-spectrogram distances computed over multiple resolutions ($80$ mel bands and hop lengths of $\approx$ $10$, $25$, and $50$ ms) as a proxy for a phase-insensitive perceptual metric. . (III) The Kernel Audio Distance (KAD) [chungKADNoMore2025] between the embedding distributions of the original and reconstructed tracks [^4] Embedding distribution metrics have become a standard way of evaluating generative models, so we include it for completeness. KAD is computed using half of the reconstructed tracks, using the other half as the reference set. We use the kadtk library with the LAION-CLAP model..
Table 2 summarizes the results. Our retrained baseline (M2L) already improves on the public M2L-Pub weights [pasiniMusic2LatentConsistencyAutoencoders2024], likely due to our distinct training data, batch size and augmentations. The linearized model (Lin-CAE) remains notably comparable on MSS and improves further on SNR, indicating that our method does not degrade reconstruction fidelity. Compared to SA-VAE, Lin-CAE is slightly worse, a similar finding to [pasiniMusic2LatentConsistencyAutoencoders2024, pasiniMusic2Latent2AudioCompression2025]. We note that SA-VAE uses different data and a two-stage training procedure that optimizes both reconstruction and adversarial losses, while Lin-CAE is trained with a single denoising objective. KAD scores for Lin-CAE are better than our retrained M2L-Pub model and SA-VAE , but are worse than the public M2L weights (possibly reflecting the different training data distribution).
@@ 4.4 Additivity
Additivity is evaluated on four source mixtures from the Musdb18-HQ test set [rafiiMUSDB18HQUncompressedVersion2019], where a mixture is : $\text{mix}=\sum^{4}_{i=1}s_{i}$. $s_{i}$ correspond to either vocals, bass, drum or other. We measure Encoder Additivity Error as the deviation of a mixture’s latent from the sum of its source latents, calculated as the relative $L_{2}$ error: $\|\operatorname{Enc}_{\theta}(\text{mix})-\sum_{i}\operatorname{Enc}_{\theta}(s_{i})\|_{2}/\|\operatorname{Enc}_{\theta}(\text{mix})\|_{2}$. For Decoder Additivity, we test the ability to reconstruct a mixture from the sum of its source latents by SNR/MSS between the decoded sum of latents, $\operatorname{Dec}_{\theta}(\sum_{i}\operatorname{Enc}_{\theta}(s_{i}))$, and the autoencoded mixture, $\operatorname{Dec}_{\theta}(\operatorname{Enc}_{\theta}(\text{mix}))$.
@@ 4.5 Music source separation via latent arithmetic
We perform source separation via latent arithmetic on the Musdb18-HQ test set. A source $\hat{s}_{i}$ is estimated by subtracting the latent representation of its accompaniment from the latent of the full mixture: $\hat{s}_{i}=\operatorname{Dec}_{\theta}(\operatorname{Enc}_{\theta}(\text{mix})-\operatorname{Enc}_{\theta}(\text{mix}-s_{i}))$. Figure 3 illustrates this process. The quality of the separated source is evaluated against the autoencoded ground-truth source, $\operatorname{Dec}_{\theta}(\operatorname{Enc}_{\theta}(s_{i}))$, using Scale-invariant Signal to Distortion Ratio (SI-SDR) [lerouxSDRHalfbakedWell2019] and MSS.
Table 1 shows that Lin-CAE significantly outperforms the baselines in latent arithmetic/oracle source separation tasks, achieving higher SI-SDR and lower MSS scores by a large margin for every instrument. The model was not trained to recover the exact phase, which is why the relatively low SDR values are observed for both reconstruction and separation. Notably, however, both separation metrics are of the same order of magnitude as the reconstruction values for the mix, which can be seen as an upper bound for the model. That is not the case, however, for any of the baselines, with the least significant gap for SA-VAE. We note that during training, we have only shown positive gains to the model, while during separation, we are effectively applying negative gains when subtracting latents. This further indicates the robustness of the learned linear structure.
@@ 5 Conclusion
We introduce a simple method to induce an approximately linear latent space in audio autoencoders during training. Using implicit conditioning via data augmentation, our approach enforces homogeneity and additivity (properties of a linear map) without modifying the architecture or objective. Experiments with Consistency Autoencoders show that linearization preserves high-quality reconstruction and can be used for source separation via latent arithmetic.
Our approach enables more interpretable and controllable audio generation. Future work could extend this method to improved CAE architectures [pasiniMusic2Latent2AudioCompression2025] and downstream tasks such as generative source separation [marianiMultisourceDiffusionModels2024]. Direct, low-level manipulations in compressed space promise new possibilities for audio editing and efficient signal processing.